solve for x and y if x squared minus y squared equals 901.
Answer by Alan3354(69443)

solve for x and y if x squared minus y squared equals 901.
--------------
There are an infinite # of solutions.
---
x^2 - y^2 = 901
y^2 = x^2 - 901
----
Pick any value for x and find y.
eg:
x = 50
y = sqrt(1599)
x = 100
y = sqrt(9099)
etc